Which of the following is represented by a parabola with vertex at the origin and focus (-3, 0)?
A. x= 1/12 y^2
B. y= -1/12 x^2
C. y= 1/12 x^2
D. x= -1/12 y^2

Answer by lwsshak3(11628)

Which of the following is represented by a parabola with vertex at the origin and focus (-3, 0)?
A. x= 1/12 y^2
B. y= -1/12 x^2
C. y= 1/12 x^2
D. x= -1/12 y^2
***
D. x= -1/12 y^2
y^2=-12x
This is an equation of a parabola that opens leftward:
Its basic equation: (y-k)^2=4p(x-h), (h,k)=coordinates of vertex
For given equation:
vertex: (0, 0)
axis of symmetry: y=0
4p=12
p=3
focus: (-3, 0)(p-distance left of vertex on the axis of symmetry)
Answer is D